Heat is added to 200 kg of ice at 0 ºC until dry steam at 100 ºC is obtained.

The specific heat capacity of water is 4200 J kg‑1 K‑1, the specific latent heat of steam is 2260 kJ kg-1 and the specific latent heat of ice is 335 kJ kg-1.

Determine the total amount of heat energy (in MJ) added to the ice.


1
Expert's answer
2021-06-29T08:22:10-0400

Gives

M=200kg

"s_w=4200jkg^{-1}K^{-1},"

Letent heat "L=2260kj\/kg"

"s_{ice}=335kJ\/kg"

0° ice to 0° water

"Q=mL_w"


"Q_1=200\times4200=840000KJ"

0° water to 100°water

"Q=ms_w\u2206T"


"Q_2=200\times335 \times(100-0)=670000KJ"

100° water to 100steam


"Q_3=m_{steam}L=200\times2260=452000KJ"

"Q_t=Q_1+Q_2+Q_3=840000+670000+452000=1.962MJ"